classes A and B have 35 students each. if 7 girls shift from A to B,then the number of girls in classes would interchange. If 4 girls shift from class B to A,then the number of girls in class A would be twice the original number of girls in class B.what is the number of boys in class A and B?
Found 2 solutions by sunny1992, tenkun:
Answer by sunny1992(35)

G1 + B1 = 35 ==> class A
G2 + B2 = 35 ==> class B
G1-7=G2 ==> eq 1
G2+7=G1
G2-4
G1+4=2G2 ==> eq 2

now taking difference of eq 1 and 2
G1-G2=7
G1-2G2=-4
___________
0+g2=11
G2=11 ==> num of girls in Class B
taking eq 1 and finding number of girls in class A
g1-11=7
g1=18 ===> num of girls in class A

Answer by tenkun(12)

let the no of girls in class A= x
the no of girls in class B= y
7 girls shifted from A to B
after shifting in class A= x-7
in class B= y+7
given condition is no of girls are interchanged
so in class A - x-7 = y
in class B - y+7 = x
this gives equation x=y+7 ---------(1)
now after shifting 4 girls from B to A
in class A = x+4
in class B =y-4
given condition is the number of girls in class A would be twice the original number of girls in class B
x+4= 2*y --------(2)
substitute (1) in (2)
y+7+4 = 2*y
2*y - y = 11
y=11
x= 11+7 =18
in class A girls = 18
boys = 35-18=17
in class B girls = 11
boys = 35-11=24
HENCE SOLVED